What is the coefficient of x^4 in y = cosh(x)






1
Expert's answer
2022-02-01T07:07:25-0500

The coefficient of "x^4" is given by "\frac{(\cosh x)^{(4)}(0)}{4!}", where "(4)" means the fourth derivative of "(\cosh x)".

We have the following relations :

"(\cosh x)'=(\frac{e^x+e^{-x}}{2})' = \frac{e^x-e^{-x}}{2} =\sinh x"

"(\sinh x)' = (\frac{e^x-e^{-x}}{2})' = \frac{e^x+e^{-x}}{2}=\cosh x"

Therefore we deduce that "\cosh x = (\cosh x)'' = (\cosh x)^{(4)}". This gives us "\frac{(\cosh x)^{(4)}(0)}{4!}=\frac{1}{4!}" as "\cosh (0) = 1". We could also arrive to the same result just using the fact that the coefficient of "x^4" in "\cosh x" is (coefficient of "x^4" in "e^x" + coefficient of "x^4" in "e^{-x}")/2.